### v2

# S2malloc: Statistically Secure Allocator for Use-After-Free Protection And More

Abstract
Attacks on heap memory, encompassing memory overflow, double and invalid free, use-after-free (UAF), and various heap spraying techniques are ever-increasing. Existing entropy-based secure memory allocators provide statistical defenses against virtually all of these attack vectors. Although they claim protections against UAF attacks, their designs are not tailored to detect (failed) attempts. Consequently, to beat this entropy-based protection, an attacker can simply launch the same attack repeatedly with the potential use of heap spraying to further improve their chance of success.
We introduce S2malloc, aiming to enhance UAF-attempt detection without compromising other security guarantees or introducing significant performance overhead. To achieve this, we use three innovative constructs in secure allocator design: free block canaries (FBC) to detect UAF attempts, random in-block offset (RIO) to stop the attacker from accurately overwriting the victim object, and random bag layout (RBL) to impede attackers from estimating the block size based on its address.
We show that (a) by reserving 25% of the object size for the RIO offset, an 8-byte canary offers a 69% protection rate if the attacker reuses the same pointer and 96% protection rate if the attacker does not, against UAF exploitation attempts targeting a 64 bytes object, with equal or higher security guarantees against all other attacks; and (b) S2malloc is practical, with only a 2.8% run-time overhead on PARSEC and an 11.5% overhead on SPEC. Compared to state-of-the-art entropy-based allocators, S2malloc improves UAF-protection without incurring additional performance overhead. Compared to UAF-mitigating allocators, S2malloc trades off a minuscule probability of failed protection for significantly lower overhead.
Keywords:
Secure Memory Allocator Use-After-Free.
1 Introduction
Heap-related vulnerabilities are serious and common threats that can be leveraged to launch attacks resulting in arbitrary code execution or information leakage. Heap overflow, double and invalid free, and use-after-free (UAF) are among the most common types of these vulnerabilities. According to the Common Vulnerabilities and Exposures (CVEs) report of 2022, they were ranked 19th, 11th, and 7th respectively, in terms of bug prevalence [27].
Secure memory allocators are an important defense against heap exploitations. State-of-the-art secure allocators can be generally classified into two categories: entropy-based generic memory allocators and UAF-mitigating allocators.
UAF occurs when a previously freed memory block is used to store data. It receives special attention in secure allocator design due to its prevalence and the powerful exploitation primitives (e.g. arbitrary read/write) it enables. Chromium has reported that more than one-third of their security issues are related to UAF, more prevalent than other types of memory errors combined [38]. While complete mitigation of UAF is theoretically possible by virtually never re-using a previously freed memory block, UAF-mitigating allocators incur substantial overheads (e.g., about 40% for MarkUs [3]), and their complexity leaves them vulnerable to new attacks [42] (albeit preventable).
On the other hand, entropy-based heap allocators aim to provide comprehensive protection against most, if not all, common heap vulnerabilities with simpler designs but may fail with a small probability. Specifically on UAF mitigation, entropy-based memory allocators typically use delayed free-lists [33, 24] to prevent the same memory block from being immediately re-allocated after being freed. Attackers now face a moving-target even when they manage to obtain a dangling pointer as they have less confidence in knowing when this pointer becomes valid again and/or which object it might point to. While achieving relatively low overhead on time usage, especially compared with UAF-mitigating allocators (see § 2 for details), existing entropy-based allocators still face the challenge of entropy-loss due to heap spraying, information leak, and silent failures on (potentially repeated) trials.
To overcome these challenges, we propose S2malloc, a novel heap allocator that combines the best of both types of memory allocators—high security assurance against not only UAF but also other types of heap vulnerabilities and yet with low memory and computation overhead on par with state-of-the-art entropy-based memory allocators. S2malloc achieves its promises through several innovative constructs: randomized in-slot offset (RIO), free-block canary (FBC), and random bag layout (RBL). RIO mitigates UAF attacks by allocating blocks with random offsets, obstructing the attacker from locating the target field in a data structure. FBC puts cryptographically secure canaries in free blocks to detect illegal writes, turning a failed UAF exploitation attempt into a clear signal. RBL organizes blocks of the same size range using sub-bags. Only blocks within the same memory page are guaranteed to be in the same sub-bag.
[…]
Summary.SummarySummary. We claim the following contributions: we
  • analyze current entropy-based allocators in real-world UAF attack scenarios and show that the level of protection they provide is not as strong as originally claimed (§ 3).
  • present S2malloc, a straightforward drop-in solution addressing the above weaknesses, while also protecting against all other commonly observed heap memory vulnerabilities. S2malloc does not require any special hardware, program recompilation, or elevated privileges, and works on both x86 and AARCH, thus increasing its deployability (§ 4).
  • through various real-world CVEs and benchmarks, show that S2malloc can successfully detect all attacks while incurring a 2.8% execution delay and 27% memory overhead on the popular PARSEC benchmark and 11.5% and 37% respectively on SPEC (§ 5§ 6).
The software artifact is open-sourced.
2 Background and Related Work
In this section, we give a succinct overview of common heap vulnerabilities and the most recent advancements in two defense lines: moving-target defense against generic heap exploitations and complete mitigation against use-after-free (UAF)—the current most frequently encountered type of memory error [38].
2.1 Heap vulnerabilities
Common heap vulnerabilities include 1 buffer overflow (reads/writes to an out-of-bound memory location), 2 double free (frees an already-freed object), 3 invalid free (frees an invalid pointer), and, as explained later in details, 4 use-after-free (UAF). Successful exploitation of these vulnerabilities can cause heap corruption, leading to devastating results such as denial of service (DoS), information leak, arbitrary code execution, and/or privilege escalation.
[…]
Allocator-assisted UAF.Allocator-assisted UAFAllocator-assisted UAF. Use-after-free (UAF) is a common heap error that occurs when a program unintentionally releases a heap object but continues to use the original pointer, i.e., a dangling pointer. Abusing the dangling pointer, an attacker can create powerful exploitation primitives such as arbitrary-address-write and code execution. A typical UAF exploitation involves the following steps:
  • Trigger the vulnerable free() function to release a heap object $A$, turning the pointer $P$ that points to the freed object into a dangling pointer.
  • Look for a victim heap object $B$ having ideally the same size as the released object. Victims containing interesting data such as pointers or length are usually preferred.
  • Wait for the program to allocate a new $B$ object that could refill the memory space initially belonging to $A$. The attacker can now corrupt the victim object $B$ with pointer $P$. This primitive can be used to either hijack the control flow or corrupt sensitive data such as uid or gid.
2.2 Entropy-based allocators
Entropy-based heap allocators strive to provide protections against almost all of the above-mentioned heap exploitations by minimizing the attacker’s success rate. Ideally, the success rate should be low enough to deter attackers from even trying to attack the system. However, practical implementations face limitations in terms of memory and computation resources, as demonstrated in Guarder [33].
[…]
BIBOP.BIBOPBIBOP. State-of-the-art entropy-based allocators typically build on the Big Bag of Pages (BIBOP) [17] memory management mechanism with various security enhancements. BIBOP-style allocators classify allocation sizes as classes. For each size class, one or several continuous pages are treated as a bag, and all allocations of the same size class will be assigned to a corresponding bag. Each bag is split into several slots preemptively and each allocated object will take one of them. The status of each slot is monitored by a bitmap and can be used to defend against double or invalid frees. Bag metadata is stored separately to avoid metadata-based attacks [32], and UAF only occurs within the same size class.
[…]
Extended secure features.Extended secure featuresExtended secure features. Prior works have introduced a diverse set of security enhancements over the basic BIBOP-style allocation, including:
  • Guard page: is a single-page virtual memory block not mapped to the physical memory. Therefore, any attempt to dereference an address in a guard page triggers a segmentation fault. Guard pages could be strategically placed after each bag or randomly within bags to protect against overflows or random accesses.
  • Heap canary: is a small object set to a secure value and placed at the end of each slot. Heap overflow can be detected if the canary value is modified.
  • Random allocation: guarantees that slots within each bag are not allocated sequentially (i.e., linear allocations). Instead, each allocated slot is randomly chosen from at least $r$ free slots. More slots will be requested if there are not enough free slots to satisfy this requirement. Intuitively, the entropy (reflected by $r$) marks a trade-off between security and performance overhead.
[…]
Evolution history.Evolution historyEvolution history. While many entropy-based allocators have been proposed, we introduce three representative works that have contributed to advancing the field.
DieHarder [29] is one of the earliest entropy-based secure allocators that adopts the BIBOP-style memory management and provides statistical protections against heap exploitations. Despite incorporating all the security features mentioned above, DieHarder has several issues compared to more recent works. These include unstable allocation entropy, predictable guard pages, and a significant impact on the execution time of the hardened program.
Guarder [33] offers multiple improvements over DieHarder, such as using a linked-list-based free block management algorithm instead of bitmap traversal, providing stable allocation entropy, and offering in-bag tunable random guard pages. Guarder significantly reduces the run-time overhead to less than 3%, making it suitable for production systems.
SlimGuard [24] further extends Guarder by reducing its memory overhead. It divides size classes into 16-byte increments instead of powers of two. However, the current implementation of SlimGuard has two security compromises: it 1) allocates blocks sequentially in the free-list, violating the claim of random allocation. 2) reuses freed blocks to store metadata, possibly due to an implementation flaw, violating the argued metadata separation practice.
2.3 UAF-mitigating allocators
UAF-mitigating allocators specialize in UAF-protection only and can be broadly categorized into the following types:
  • A
    Pointer validation on dereference [40, 14, 43]: i.e., checks whether a pointer is valid and safe for a read/write operation upon dereference.
  • B
    Pointer invalidation on free [16, 10, 31, 15, 42]: i.e., whenever free is called on a pointer, it revokes its capability to access the associated virtual address.
  • C
    Memory sweeping [20, 3, 13]: i.e., checks all pointers stored in memory and either actively removes all dangling pointers or reuses a freed memory chunk with the assurance that no dangling pointer to this freed chunk exists.
  • D
    Context-based allocation [4, 19, 12]: i.e., permits the re-allocation of freed memory chunks only to objects allocated in the same “context.”
Not all UAF-mitigating allocators are drop-in replacements of the system memory allocator, as some of them require recompilation of the protected programs [4, 19, 41], special hardware [15], or kernel modifications [16].
[…]
Security guarantee.Security guaranteeSecurity guarantee. UAF-mitigating allocators in categories A, B, and C can eliminate all UAF exploits (hence complete mitigation), although some of them might incur large overheads or require customized kernel or hardware. Context-based allocators (category D), first proposed in [12] as a heap-safety property, typically run faster but offer incomplete protection. For example, Cling [4] defines the allocation context based on the two innermost return addresses on the call stack when malloc is invoked. Suppose two memory allocations occur in the same context; in this case, the object allocated the second time may occupy the same memory chunk allocated the first time (if the first object is freed). TypeAfterType [19] defines the allocation context based on object types (e.g., the type int passed in malloc(sizeof(int)), and objects may be reallocated on freed memory chunks originating from the same allocation context. In summary, UAF is still possible in context-based allocators.
However, unlike entropy-based allocators (including S2malloc), which also mitigate UAF imperfectly, the protection from context-based allocators is deterministic. Specifically, if the dangling pointer and the target object are allocated in different contexts, there is zero chance of success in corrupting the target object, regardless of the attackers’ strategies. On the other hand, context-based allocators provide no protection if the dangling pointer and the target object fall in the same context. Entropy-based allocators provide probabilistic protection in both cases. We will discuss the implications via CVEs in § 5.5.
[…]
Porting for UAF-write defense only.Porting for UAF-write defense onlyPorting for UAF-write defense only. Allocators that instrument runtime checks for pointer validity (i.e., category A) can opt to trade protections of UAF-read attacks for performance by checking the pointer validity on write accesses only. In fact, as reported in [34], only 5% to 25% of memory accesses are write accesses in the SPEC 2017 benchmark, indicating a potential reduction in overhead for category-A allocators. On the other hand, for allocators in categories B, C, and D, splitting UAF-read and UAF-write protection is inherently hard as their designs do not differentiate between read and write accesses.
3 Motivation
While effective in defending against various heap exploits, entropy-based allocators are not ideally suited to protect against UAF attacks. This partially motivates the stream of research on UAF-mitigating allocators presented in § 2.3. Specifically, there exist easy-to-identify blind spots that drastically reduce the efficacy of defending UAF using random allocation and increase attackers’ confidence on launching a successful attack without trying to predict the next allocation slot. In this section, we first present the adversary model followed by a discussion on the limitations of existing entropy-based allocators, which motivates the development of S2malloc.
3.1 Adversary model
We assume that the attacker can analyze the source code and binary executable to determine the implementation details of the victim program, including vulnerabilities and other relevant information such as the size and layout of critical data structures. We also assume that the attacker can identify when a victim object is allocated or de-allocated.
We do, however, assume that the underlying kernel and hardware are trusted and an attacker cannot utilize a data leakage channel, like /proc/\$pid/maps, to discover the location of the heap allocator’s metadata. The attacker cannot compromise the random number generator nor can they take control of the heap allocator. Exploiting bugs of the allocator itself is out of scope. These assumptions are similar to that of other entropy-based allocators [33, 24].
Additionally, we allow attackers to use any existing heap feng-shui [35] technique (e.g., heap spray) to prepare or manipulate the layout of heap to facilitate UAF-exploits. And attackers can retry an exploit as long as previous attempts fail silently. These assumptions make our adversary model stronger than those assumed in entropy-based allocators and on-par with the adversary models in UAF-mitigating allocators [3, 42].
3.2 Challenge 1: entropy loss
Entropy-based allocators thwart UAF by avoiding instant memory reuse. However, if 1) the attacker could continue to retry the attack when the previous trial fails, or 2) the heap memory can be spoofed with either dangling pointers or victim objects, it is guaranteed that the attack would eventually succeed even without the victim’s notice.
Figure 1 is an example, abstracted from mRuby issue 4001 [11], a UAF vulnerability in the Ruby compiler. The function mrbioinitializecopy is called when opening a file. It first frees the existing data pointer of the copy object (DATAPTR(copy)) (line 7) and allocates new memory to it (line 9 and 11). If an invalid argument is passed, calling iogetopenfptr would throw an exception (line 10), making DATAPTR(copy) a dangling pointer to the freed object.
Using this vulnerability, the attacker can allocate a string object to take the freed memory space. The attacker can then close the file in Ruby, which will set the first word of the pointed memory to INF. If this memory is taken by a string object, its length will be overwritten to INF that allows arbitrary memory read and write.
1mrb_io_initialize_copy(mrb_state *mrb, mrb_value copy) {
2  mrb_value orig;
3  struct mrb_io *fptr_copy, *fptr_orig;
4
5  fptr_copy = (struct mrb_io *)DATA_PTR(copy);
6  if (fptr_copy != NULL) {
7    mrb_free(mrb, fptr_copy);
8  }
9  fptr_copy = (struct mrb_io *)mrb_io_alloc(mrb);
10  fptr_orig = io_get_open_fptr(mrb, orig);
11  DATA_PTR(copy) = fptr_copy;
12}
Figure 1: Example UAF attack based on mRuby issue 4001 [11]
In the above scenario, random (i.e., non-sequential) allocation or delayed free-lists available in existing entropy-based allocators [29, 33, 24] merely increases the attack difficulty: as long as the attacker can wait, the memory chunk referred to by the dangling pointer will eventually be re-allocated, allowing the UAF exploit to proceed after some delay.
Furthermore, the entropy diminishes if an attacker is allowed to repeat the same attack without penalty (e.g., when a failed attempt does not crash the target program or trigger attention). Similarly, if attackers have the ability of spraying the heap with either dangling pointers of victim objects, the probability of success increases significantly.
This motivates us to design an allocator in a way that 1) actively searches for UAF attempts and raises signals if the evidence is found; and 2) stops the attacker from locating critical information in memory even if the attacker manages to obtain a dangling pointer. In this example, we could prevent the attacker from being able to deterministically locate the string length even if the attacker manages to obtain a dangling pointer to a string object originally pointed to by DATA_PTR(copy). In addition, any attempts of writing to unallocated memory will be detected with a high probability. If an attacker attempts to spray the heap with arbitrary write to increase success rates, we can raise a signal on or even before the UAF actually happens.
3.3 Challenge 2: information leak
Existing entropy-based allocators [33, 24] create a huge memory pool for each block size range, resulting in the leakage of block size via their address, possibly revealing the victim software’s internal state to the attacker. For example, each BIBOP bag is assigned an 8 Gigabytes virtual memory pool in SlimGuard [24]. Any objects belonging to this bag will be allocated from this pool. As a result, obtaining a known sized block will be sufficient enough to infer the size of any blocks sharing the upper 32-bit address. Further, block addresses in Guarder [33] are aligned by their size. Block size would possibly be inferred just based on its address. For example, a block with an address value ending with 0x10 is of size 1 to 16 bytes, and a block with an address value ending with 0x300 is highly likely of size 129 to 256 bytes.
We mitigate this threat by dividing bags into sub-bags, limiting the size leakage only if the attacker-controlled block resides in the same sub-bag as the victim block. Furthermore, we assign random guard pages within sub-bags to make the sub-bag boundaries unpredictable.
4 Design and Implementation
Now we explain the design and implementation of S2malloc and how it thwarts the types of heap attacks in § 2.1.
4.1 Architectural overview
At its core, S2malloc adopts BIBOP to manage memory blocks. An overview of S2malloc is illustrated in Figure 2. S2malloc maintains a per-thread metadata ( A), stored in a memory chunk requested directly from the kernel. Huge blocks are obtained or released from the OS directly ( B), and are stored using a linked list. Small objects are maintained using bags, claiming memory indirectly from a segregated memory pool ( C). Each regular bag maintains the metadata of blocks of a size range, including the number of free slots and a list of sub bags. We take sub bag as the basic unit of a group of slots  (§ 4.3).
The data field points to a memory chunk requested from the memory pool to store the objects allocated to this sub-bag. The bitmap indicates whether the current slot is taken or not. If the current slot is taken, the corresponding offset table cell stores an offset indicating where the data stored in the bag starts (§ 4.2). Otherwise, the slot is free and the offset table stores the location of the FBC (§ 4.5).
When a free() call is received by S2malloc (step 1 $\rightarrow$ 2), S2malloc checks whether: (1) the bitmap indicates the current block is taken; (2) the offset stored in the offset table matches with the freed pointer address; and (3) the canary value is not modified. If all checks pass, the current block will be freed: the bitmap cell will be set to free and an FBC will be put at a random location in the current block. The offset table will be updated to store the location of the FBC.
When a malloc() call comes, (step 2 $\rightarrow$ 3), S2malloc randomly selects one free block of the corresponding size and checks the FBC of current and nearby free blocks. A random offset will be generated indicating where the data starts within the current block, and the offset table will be updated accordingly. The heap canary will be set after the last data byte of the current block, and the bitmap will be updated. In this example, we assume that each block stores at most 7 bytes of data. The heap canary is set to the 9th byte initially as the offset is one and is then set to the 10th after reallocation as the offset is changed to two.
4.2 Randomized in-slot offset (RIO)
Figure 2: Overview of S2malloc with an example of free and malloc. A, B, and C show three S2malloc segments, stored in segregated memory. 1, 2, and 3 show how an allocated bag slot is freed and then allocated.
In all existing secure memory allocators, allocated objects store their data from the first byte of the allocated slot. Alternatively, we propose that the object will be stored with a random offset $p$, and the first $p$ bytes of the slot will be left empty. After the slot is freed and allocated to another object, the offset $p$ will be re-computed. Thus, the relative offset between these two objects cannot be accurately predicted and the attacker cannot accurately re-use a freed pointer and arbitrarily read or write the target memory.
We define $k$ to be the RIO entropy. For each bag with blocks of $b$ bytes, $e=b/k$ bytes are not used to take data, and each block can take at most $b-e$ bytes of data, guarantee minimal in-slot offset entropy. We refer to these extra bytes as entropy bytes. Suppose this block is malloc-ed with an object of $s$ bytes ($s<b-e$). Before this block is allocated, $p\in(0,b-s)$ is computed to decide the starting byte of the data object. $p$ is 16-byte aligned following the minimum default alignment of GNU C implementation [39], and to be compatible with special data structures, such as atomic objects that need to be stored to align with the registers and cache lines. The offset of each block will be stored separately in an offset table. The minimal entropy $e$ increases as $b$ becomes larger to avoid introducing high memory overhead for small objects and to provide stronger protections for large objects, observing the fact that larger objects have more complicated structures and are more likely to be targeted.
We adopt the Permuted Congruential Generator (PCG) [30] algorithm following the design of  [24] to generate all random numbers. With negligible execution time – much faster than existing random number generators, such as LCG and Unix XorShift, PCG generates hard-to-predict numbers. While it does not provide cryptographic security, the only known attack towards PCG requires three consecutive PCG outputs to recover the seed [7]. However, in S2malloc, none of the generated numbers is accessible to the user and can only be obtained by sweeping the memory. As a result, PCG is sufficient enough to provide the required level of entropy.
4.3 Random bag layout (RBL)
As with existing secure allocators, S2malloc employs BIBOP-style management for small-size blocks. Blocks larger than 64 kilobytes are mapped and unmapped from the OS directly, and are managed using a linked list. Blocks smaller than 64 kilobytes are further classified as small, medium, and large blocks to decrease the number of size classes. Small bags contain blocks smaller or equal to 1 kilobyte, and a bag is created every 16 bytes (16 bytes granularity). For example, the first small bag takes blocks smaller than 16 bytes, and the second small bag takes blocks of (16, 32] bytes (without taking RIO into consideration). Medium bags contain blocks within the range of 1 kilobyte and 8 kilobytes with the granularity of 512 bytes; large bags contain blocks within the range of 8 kilobytes and 64 kilobytes with the granularity of 4 kilobytes. In total, S2malloc has 64 small bags, 14 medium bags, and 14 large bags.
S2malloc obfuscates the virtual memory allocation and stops linking block sizes to their addresses. Instead of allocating a dedicated virtual memory pool for each bag, (as shown in prior works [33, 24]), we create a single virtual memory address pool for all bags. We further divide each bag into sub-bags each containing 256 slots. Each bag creates new sub-bags upon need, and a newly created sub-bag would request corresponding memory from the pool. We use a bump pointer to track the available memory in the pool and linearly allocate pool memory to sub-bags.
S2malloc randomly places guard pages within sub-bags to thwart overflow, spraying, and random pointer access. If a sub-bag is randomly allocated with a guard page, one of its pages will be unmapped (protected) randomly using the mprotect system call. Any slots within this protected page will be marked as allocated in the bitmap to avoid allocating them to the program. Any accesses to these slots are thus invalid and result in a segmentation fault. The tunable guard page rate can be configured in an environment variable. We note that the memory pool allocation is not deterministic and cannot be predicted due to the random guard pages.
S2malloc guarantees that the block size leakage occurs only if the known block and the victim block reside on the same memory page: adjacent blocks may not be within the same sub-bag, and RIO guarantees that blocks start at addresses that cannot be deterministically predicted. On the contrary, two blocks are highly likely of the same size range in SlimGuard if their address difference is smaller than 8GB.
Secure Guard Rand. Segre. Heap Ptr UAF UAF Overheads
Allocators Pages Alloc. MD Canary Inval. Mitigate Detect Memory Runtime
DieHarder Y Y Y N N N 21.3% 2.1%
Guarde Y Y Y Y N N 58.1% 2.4%
SlimGuard Y N N Y N N 22.5% 4.4%
S2malloc Y Y Y Y N High prob. 26.8% 2.8%
MarkUs N N N N Y N 13.0% 42.9%
FFmalloc N N N N Y Y 50.5% 33.1%
Table 1: Overview of existing secure memory allocators and S2malloc to illustrate how S2malloc fills the gap (MD: metadata, ◐: One-shot attack only, ●: Repeated/spray attacks). Overheads are measured on PARSEC [6], detailed in § 6. Overheads of MarkUs and FFmalloc (numbers marked with $\ast$) are reported in [42] instead of measured by us.
4.4 Hardening heap canaries
Canary is a small data block put after the allocated memory to detect overflow, initially introduced in StackGuard [8] to protect the stack. Canary has now been adopted to protect the heap [28]. At the time a memory slot is being allocated, the canary will be set to a specific value. This value will be checked at the time this slot is freed, and memory overflow will be detected if the canary value changes.
However, in previous designs, this value is set to be either globally identical [33] or is binded with slots [24], and could be trivially broken by a knowledgeable attacker. We follow the design of previous works to use the secure MAC of the memory address as the canary [21, 26]. Specifically, we take the CMAC-AES-128 encrypted block address as the canary implemented using AES-NI [18] (on x86 CPUs) or Neon [5] (on AARCH CPUs) to keep the canary confidential and compact. Even if the attacker learns a canary value, they can only use it to break the current object or any further objects allocated to this slot with the same RIO.
Specifically in S2malloc, we put a $\iota$-byte canary immediately after the last data-storage byte in the allocated slot, (i.e., the $(p+b-e)$th byte), and the canary will be checked upon free.
4.5 Free block canaries (FBC)
Existing entropy-based allocators defend against UAF-write attacks by statistically avoiding allocating a victim object in a block pointed to by a dangling pointer. Although a failed attack attempt only modifies a free block without causing any harm, the attack attempt is not detectable either and given the fact that the same attack can be retried, the attacker will succeed eventually.
To detect such attempts, we put a canary of length $c$ in each free block. The canary value is also computed using CMAC-AES-128. This value will be checked before the block is allocated and will be reset after it is freed. We also check the FBC of $d$ nearby blocks to improve the rate of detection. FBC guarantees that accessing a freed block is not risk-free. Its protection rate is analyzed in § 5. In Figure 2, we further illustrate how the two kinds of canaries (FBC and regular heap canary) are set and cleared when an allocated block is first freed and then allocated again.
Initially in S2malloc, we create the memory pool using the mmap system call with the ANON flag and all allocated memories are set to zero in the Linux environment [22]. We take this advantage and use the zeros as the initial FBC with the following benefits:
  • Until being accessed, an unused slot will remain unmapped to the physical memory, decreases memory overhead.
  • An unused slot is exempted from computing a secure canary value and writing to the corresponding memory field.
  • The whole slot will be checked instead of only the canary bytes, increasing the detection rate.
Treading off the computation cost of the encrypted canaries, S2malloc always zeros out the contents of small blocks and will check the whole block before being allocated to a new object, bringing both security and computation benefits.
4.6 Summary and comparison
Table 1 summarizes S2malloc and selected state-of-the-art secure heap allocators along the two defense lines that are closely related to S2malloc (background discussed in § 2.2 and § 2.3).
Being an entropy-based allocator, S2malloc is inherently closer to this line of work [29, 33, 24] with a nearly identical set of heap exploitation protection features except UAF protection. S2malloc provides a much stronger security assurance in the presence of UAF vulnerabilities. In particular, S2malloc addresses the two entropy-loss cases (discussed in § 3.2 and § 3.3) with RIO (§ 4.2) and RBL (§ 4.3), respectively, and hence, providing much higher effectiveness on UAF mitigation. In addition, S2malloc is designed to actively monitor the integrity of the heap and watch for UAF attempts, including heap spraying practices that aim to prepare the heap data and layout for UAF exploits. S2malloc achieves this through a synergy of regular heap canaries (§ 4.4) and FBC (§ 4.5).
On the other hand, Table 1 also shows a sheer contrast between entropy-based allocators and UAF-mitigating allocators. Notably, although providing a theoretically complete mitigation guarantee toward UAF, UAF-mitigating allocators [3, 42] significantly impair program efficiency and are hard to be deployed in time-sensitive use cases. In contrast, as will be presented in § 6, S2malloc incurs a significantly lower overhead that is typical for entropy-based allocators, making S2malloc practical and deployable on production systems if the end-user can tolerate a marginal chance of protection failure (less than 10% in the default setting of S2malloc, discussed in § 5).
5 Security Evaluation
(a) S1 protection rate
(b) S1 attack success rate
(c) S2 protection rate
(d) S2 attack success rate
(e) S1-spray protection rate
(f) S1-spray attack success rate
(g) S2-spray protection rate
(h) S2-spray attack success rate
Figure 3: Parameterized security evaluation (x-axis: logarithmic block size/byte, y-axis: protection/attack success rate/%).
In this section, we show the robustness of S2malloc towards UAF exploitations in different scenarios. We first present the results from our formal modeling and then show how S2malloc mitigates real-world UAF attacks.
5.1 Formal analysis
To mathematically model how S2malloc provides defense against UAF, we make the following assumptions for the attacker and target program (which are consistent with our adversary model in § 3.1):
  • 1
    The goal of the attacker is to modify the victim field, a sensitive field (e.g., a function pointer or an isadmin flag) in a specific type of object, a.k.a., a victim object, via memory writes over a dangling pointer (i.e., UAF-writes).
  • 2
    The attacker can obtain a dangling pointer through a bug in the program at any point of time during execution.
  • 3
    The program repetitively allocates and frees the type of objects targeted by the attacker (i.e., victim objects) during its execution. However, we do not assume that each victim object is freed before the next victim is allocated.
  • 4
    The attacker can either indirectly monitor or directly control the allocations of victim objects, i.e., the attacker knows when a victim object is allocated, but does not know the address of the allocation.
  • 5
    Any memory writes through the dangling pointer is conducted after the victim object is allocated.
  • 6
    If the intended sensitive field of a victim object is overridden, the attack succeeds; otherwise, the program continues to execute, allowing the attacker to repeat the exploitation effort unless detected by S2malloc (condition 7).
  • 7
    S2malloc checks FBCs on each heap allocation and detects the attack if any FBC is modified.
To simplify the illustration, we assume that the above execution logic is the only code logic that involves heap management. In real-world settings, attackers usually have an even lower success rate as memory slots can be allocated to other objects, which gives S2malloc more chances to check FBCs and detect UAF attempts.
In addition, in reality, there will be objects that are not of the victim type but bear similar sizes and hence, will also compete for the same block with victim objects. Such objects are called irrelevant objects and their prevalence and longevity at run time affects a realistic attacker’s strategy in attacking S2malloc. Below, we introduce two simple and yet realistic strategies a reasonable attacker might consider:
[…]
S1: Repeat UAF-writes through the same dangling pointer.S1: Repeat UAF-writes through the same dangling pointerS1: Repeat UAF-writes through the same dangling pointer. If an attacker is confident that the memory block pointed to by the dangling pointer is not taken by a long-living irrelevant object, the attacker may prefer to keep re-using that dangling pointer for future attack attempts and hope that a victim object is allocated in that block.
[…]
S2: Repeat UAF-writes with freshly obtained dangling pointers.S2: Repeat UAF-writes with freshly obtained dangling pointersS2: Repeat UAF-writes with freshly obtained dangling pointers. If an attacker worries that irrelevant objects can be long-lasting (i.e., holding a block that may be allocated to victim objects), the attacker may prefer to use a fresh dangling pointer per each attempt. However, we note that this is only an attack strategy and we assume that there is no heap allocation other than victim objects.
For each of the scenarios above, we discuss the rates of whether the attacker adopts the heap spray technique or not separately. We introduce a total of four scenarios. We refer readers to Appendix 0.E for a detailed explanation of them.
5.2 Parameterized protection rates
In this section, we illustrate how the protection and attack success rates vary with different parameter configurations using the two attack strategies. Assuming the victim field is a pointer of 8 bytes, The set of tunable parameters include 1 block size, 2 FBC length, 3 RIO entropy, 4 break on free, and 5 number of free blocks with FBCs checked.
For FBC-checking, we take 0 and 4 nearby blocks as a comparison to the default setting (2 nearby blocks). For FBC length, we take 4 and 12 bytes as a comparison with the default (8 bytes). For random offset entropy, we reserve 50% of block size for RIO compared to the default (25%). We also evaluate the influence on the two rates with break-on-free.
We provide estimates of the rates by assuming that an attacker re-tries an attack for 500 times at max even if the attack is still not detected by the defender. We also assume that overwriting the heap canary does not trigger any alarm. We analyze the protection and attack success rates of blocks with sizes ranging from 16 bytes to 64K bytes as larger blocks are resistant to UAF attacks. All these simplifications favor the attacker.
3(a) shows how the protection rate changes with a different set of tunable parameters while 3(b) shows how the attack success rate changes using strategy S1. We observe that adopting a more secure setting increases the protection rate and decreases the attack success rates for both small and large blocks. However, as the block size increases, both rates decrease as it is less likely to overwrite either the target field or FBC. Both the attacker and the defender are likely not to succeed, and the rates are thus not summed up to one (in fact, not even close to one).
3(c) shows how the protection rate changes with the same set of tunable parameters while 3(d) shows how the attack success rate changes using strategy S2. Similarly, we observe that a more secure setting increases the protection rate and decreases the attack success rates. However, the attacker would leave numbers of overwritten FBC in the memory pool in this strategy making the protection rate higher.
Note that other entropy-based allocators provide zero protection as they do not detect failed UAF-write attempts.
5.3 Protection rates with heap spray
Heap spraying boosts attackers’ chance of overwriting a sensitive field in a victim object and hence, increases the attack success rate and decreases the protection rate. (see detailed strategies and formal analysis in § 0.E.4 and § 0.E.5). We use the default settings of S2malloc and evaluate how both attack and defense rates change of strategies (S1 and S2) with different numbers of sprayed objects and object sizes. 3(e) and 3(f) show the protection and attack success rates of reusing the same pointer (strategy S1-spray). 3(g) and 3(h) show the protection and attack success rates using a fresh pointer (strategy S2-spray).
Spraying the heap with target variables diminishes the random allocation entropy, thus decreases the protection rate and increases the attack success rate as FBC will not be overwritten if the pointed block is allocated thus making the attack not detectable. RIO entropy is not influenced by heap spray. An effective protection can be achieved by adopting a more secure configuration (e.g., checking more nearby FBCs, larger RIO or block entropy), with marginal performance degradation (see § 6.5) to make the spraying less effective.
5.4 Illustrate the protection rates
We take the mRuby issue 4001 (shown in Figure 1) as an example and show how its protection rate is computed. The size of object mrbio is 16 bytes, In each run, the attacker’s goal is to overwrite 4 bytes of it. With the default settings ($r=256$, $s=32$ to store a 16-byte object), the attack success rate of each trial is approximately $0.002$, and the probability of overwriting FBC in a free block is approximately $0.16$. In Table 2, we show how the rates change as the number of attack rounds goes up. The attack is 64% likely to be detected if the attacker adopts the first attack strategy and 95% likely to be detected using the second strategy after running it 500 times.
Attack Strategy 1 Attack Strategy 2
round 1 5 10 50 100 500 1 5 10 50 100 500
$p_{\_}{protection}$ 1.4% 4.1% 7.4% 28% 43% 64% 0.8% 12% 37% 95% 95% 95%
$p_{\_}{attack}$ 1.2% 2.6% 4.4% 15% 24% 35% 1.2% 2.6% 4.0% 5.5% 5.5% 5.5%
Table 2: Protection and attack success rates of attack rounds in mRuby issue 4001 using the two strategies.
5.5 Defending against real-world CVEs
In this section, we take seven recent vulnerabilities and compare how S2malloc, Guarder, DieHarder and SlimGuard fare in defending against them. We select vulnerabilities based on the following criteria:
  • On the Linux platform and can be mitigated in the user space (i.e., not a Linux kernel bug);
  • Can be deterministically triggered (i.e., not racy);
  • Public exploit is available and the exploit breaks the program information integrity (i.e., not only causing DoS).
1typedef union _zvalue_value {
2    long lval;
3    double dval;
4    struct { char *val; int len; } str;
5    HashTable *ht;
6    zend_object_value obj;
7} zvalue_value;
8
9typedef struct _zval_struct {
10    zvalue_value value;
11    zend_uint refcount;
12    zend_uchar type;
13    zend_uchar is_ref;
14} zval;
Figure 4: Type definition of zval.
Vulnerability Attack pattern [33] [29] [24] S2malloc [19] [4]
CVE-2015-6831 DP $\rightarrow$ Write
CVE-2015-6834 DP $\rightarrow$ Write
CVE-2015-6835 DP $\rightarrow$ Write
CVE-2015-6835 DP $\rightarrow$ Write $\rightarrow$ sleep
CVE-2020-24346 DP $\rightarrow$ Write
Python-91153 DP $\rightarrow$ Write $\blacksquare$
mruby-4001 DP $\rightarrow$ Write $\blacksquare$ $\blacksquare$
CVE-2022-22620 DP $\rightarrow$ Read
[…]
$\bullet$
Table 3: Sumsmary of how different memory allocators defend against eight exploitation techniques on seven vulnerabilities. Vanilla BIBOP allocator and Scudo [25] are vulnerable to all attacks and behave similarly to Guarder [33] (DP: dangling pointer, ○: no defense, ◐: detect at the end of execution, ◑: defense via zero-out, ●: detect via FBC change,
[…]
$\bullet$
: non-deterministic leak (RIO), $\blacksquare$: thwart the exploitation ability).
Out of the seven vulnerabilities, six are UAF-write bugs and one (CVE-2022-22620) is a UAF-read only bug. We also found seven exploits against these vulnerabilities (two exploits for CVE-2015-6835 with different attack patterns). All CVEs except Python-91153 and mruby-4001 can cause arbitrary code execution (ACE) if properly exploited. However, a powerful attack (e.g., ACE by overriding a function pointer) can succeed when the target object is precisely allocated to a freed memory chunk that is still referred to by a dangling pointer [^1] An attacker may attack blindly, e.g., overriding a code pointer through the dangling pointer regardless of whether it points to a target object or not. This will have three consequences: 1) corrupting FBC, which may cause the attack to be detected upon future mallocs, 2) overwriting a wrong field due to RIO which may cause the program to enter a weird state (e.g., crash), 3) a successful attack. If the program can be recovered from a weird state automatically (e.g., crash resilience), the attacker can retry the same attack and eventually case 1 or case 3 will occur. However, without the probabilistic detection on UAF attempts by S2malloc, only case 3 will occur.. S2malloc can mitigate attacks by reducing the chance that a target object is referred to by a dangling pointer. Evaluation results of the eight exploits are in Table 3.
[…]
Entropy-based allocators.Entropy-based allocatorsEntropy-based allocators. S2malloc can thwart all UAF-write attacks evaluated, Guarder can detect three exploits by recognizing double-free attempts, but DieHarder and SlimGuard fail to thwart these exploits. For the UAF-read attack, S2malloc uses RIO to stop the attacker from reusing the memory chunk with accurate object alignment, causing the data read by the dangling pointer to be not sensible. DieHarder zeros out the memory after it is freed, which is effective if the attacker tries to over-read a freed block.
[…]
Context-based allocators.Context-based allocatorsContext-based allocators. While we expect context-based allocators to demonstrate strong and stable protection, some of the exploits, unfortunately, hit on certain blind spots in Cling and TypeAfterType by accident.
In the case of Cling, if both the dangling pointer and the target object (i.e., the object an attacker hope to corrupt) are allocated through the same multi-layer function call sequence, they are considered to fall under the same allocation context, causing the target object to be possibly accessed by the dangling pointer. We will illustrate this limitation through the CVE-2015-6835 case study presented later. In fact, all examined CVEs, except mruby-4001, hit this limitation of Cling. Cling mitigates mruby-4001 by limiting the attacker to target objects of type mrb_io, which prevents the attacker from creating a powerful attack primitive.
TypeAfterType can unpack malloc wrappers with an arbitrary number of layers until it finds a sizeof(T) in the function argument, and an ID i is given to each allocation site of T. The tuple (i,T) makes the allocation context, and all memory allocations through this call sequence will be allocated from a memory pool dedicated to this context. However, if the dangling pointer and the target object share the same context in an exploit, UAF can still occur. We will illustrate this limitation through the CVE-2015-6835 case study. TypeAfterType mitigates Python-91153 by limiting the target object to be a reallocated string, and mruby-4001 by limiting the target object to be an mrb_io.
[…]
Case study: CVE-2015-6835.Case study: CVE-2015-6835Case study: CVE-2015-6835. This CVE is a UAF bug in the PHP session deserializer, which reconstructs a session from a serialized string. (Figure 7). An attacker can exploit this vulnerability to control a dangling pointer to a freed zval object. This is possible as the return value (a zval pointer) phpvarunserialize is freed in its caller without noticing that the same pointer might also be stored in a global variable SESSIONVARS.
The zval type, unfortunately, is a reference-counting wrapper over nearly all other objects involved in the PHP engine (see definition of zval in Figure 4). Therefore, an attacker might corrupt any zval object that may be reallocated to the freed slot. They can simply uses the echo(..) function to dump a newly allocated zval in the freed memory.
1) Protection from entropy-based allocators. S2malloc checks FBC on every malloc(). In this exploit, when the attacker tries to use the dangling pointer in zendechohandler, its refcount field is increased, causing the FBC to be modified. This enables S2malloc to detect the UAF attempt when the corrupted slot or a nearby slot is about to be reallocated. If the refcount change does not corrupt FBC (simulated by disabling the FBC check) and this corrupted block is reallocated, S2malloc can still stop the exploit as RIO causes misalignment between the dangling pointer and new object, causing the type field to have value UNKNOWN and prevents echo printing.
Guarder and DieHarder try to mitigate this attack by random allocation: hoping the new object will not be referred to by a dangling pointer. However, our experiment shows that Guarder fails if the attacker re-runs the attack multiple times or spray the heap with victim objects. SlimGuard fails to provide protections as it always allocates the most recently freed objects to the program. It does not implement the claimed random allocation feature and does not have any other security features that could detect UAF. DieHarder zeros out the memory chunk that stops information leakage of the freed zval, but it cannot prevent an attacker to corrupt the newly allocated zval over the freed chunk.
[…]
For Cling, this malloc wrapper implies that the allocation of many zval objects will be sharing the same context (measured by the two innermost return addresses on the call stack). This leaves the dangling pointer plenty of candidate objects to refer to after several rounds of deserialization in PHP. TypeAfterType can inline malloc wrappers but the inlining stops at phpvarunserialize because it sees the sizeof(zval) argument in emalloc and hence, will allocate all zval objects originating from this malloc wrapper from the same pool. Unfortunately, the dangling pointer is also allocated this way, enabling UAF among the dangling pointer to other zval objects as well.
[…]
Summary.SummarySummary. The combination of random allocation and delayed free-list provided by previous entropy-based allocators focus on one-time attacks only, Hence, repeating the same attack is a simple yet effective solution to undermine their protection. Context-based allocators, on the other hand, might fail to detect UAF among objects allocated of the same context. While these results highlight the effectiveness of S2malloc, we note that information leakage through corrupted pointers might diminish this guarantee, which is discussed in § 7.
6 Performance Evaluation
In order to evaluate the performance and memory overhead of these allocators, we run various benchmarks trying to provide a complete understanding of their performance. We firstly run two macro benchmarks – PARSEC and SPEC (§ 6.1), and then use the mimalloc-bench and glibc micro-benchmark to evaluate the performance of running two most frequent heap memory management functions: free() and malloc() (§ 6.2). We then evaluate their performance on real-world programs using two servers: Nginx and Lighttpd, and two databases: Redis and SQLite (§ 6.3). We then discuss how multi-threading impairs each of them performance (§ 6.4). In the end, we show how different parameter values influence the performance of S2malloc (§ 6.5).
[…]
Experiment setup.Experiment setupExperiment setup. Experiments are performed on both x86 and AARCH servers for macro benchmarks. The performance of benchmarks is measured only on the x86 server. The x86 server is configured with 64-bit 160-core 2.40GHz Intel Xeon E7-8870 CPUs (x86 architecture) with 1TB system memory. We set up the AARCH server on Amazon Web Service (AWS), using the im4gn.4xlarge machine with 16 vCPU cores and 64 GB memory. On both machines, benchmarks are measured in the Docker environment with Debian 11, kernel version 5.15.0. We measure the overheads using the GNU time binary [23] and setting the LD_PRELOAD environment variable to substitute the system default allocator.
We obtain SlimGuard, Guarder, and DieHarder from their corresponding GitHub repository. We use SlimGuard with commit 81f1b0f as a later erroneous commit prevents us from using LD_PRELOAD to replace the system allocator. We use the up-to-date version of the other two memory allocators (Guarder: 9e85978, DieHarder: 640949f). In order to provide a fair result, we reduce the allocation entropy bit of Guarder to eight (same as the default value of SlimGuard and S2malloc). We also disable DieHarder from zeroing out freed blocks (this actually slightly accelerates DieHarder).
S2malloc is measured with the settings of checking two nearby blocks ($d=2$), 10% random guard page, and taking 1/4 of the block size as random offset entropy ($e=0.25b$). For blocks smaller than a memory page (4096 bytes), we zero it out and take the whole block as FBC. For blocks larger than a memory page, we set an 8-byte FBC ($c=8)$ in the corresponding blocks. We set the heap canary length to be one byte ($\iota=1$) following the design of SlimGuard and Guarder. All reported times and memory usage are normalized using the baseline (glibc) output. We use geometric averages to compute average overheads and report the means and standard deviations of five runs.
Run-time Overhead Memory Overhead
x86 AARCH x86 AARCH
SPEC PARSEC SPEC PARSEC SPEC PARSEC SPEC PARSEC
S2malloc 12% 2.8% 16% 1.8% 37% 27% 38% 28%
SlimGuard 17% 4.4% 7.7% 2.6% 57% 23% 57% 24%
DieHarder 31% 2.1% - 2.5% 59% 21% - 21%
Guarder 3.5% 2.4% - - 56% 58% - -
Table 4: Normalized overheads on SPEC and PARSEC benchmarks.
6.1 Macro benchmarks
[…]
PARSEC.PARSECPARSEC. We first evaluate the performance of S2malloc using the PARSEC [6] benchmark. We exclude three network tests (netdedup, netferret, and netstreamcluster) and one test (x264) that fails to compile in the baseline scenario, and only report the result of the rest 12 benchmarks. Additionally, we exclude “raytrace” from execution for the AARCH sever as it cannot compile. We refer to each PARSEC test using the first three letters of its name.
[…]
SPEC CPU2017.SPEC CPU2017SPEC CPU2017. We use SPEC CPU2017 [36] version 1.1.9. We report the results of 12 C/C++ only tests in both “Integer” and “Floating Point” test suites with the default OpenMP settings of four parallel threads. All reported SPEC overheads are “reportable” following the SPEC documentation [36].
[…]
Results.ResultsResults. We measure the performance of S2malloc and three other entropy-based allocators (SlimGuard, Dieharder and Guarder) on both benchmarks with the x86 machine. On the AARCH machine, we exclude Guarder from analysis, noticing that Guarder relies on AES-NI [18], an Intel CPU extension, not supported on AARCH machines. We report the average and standard deviations of overheads in Table 4. Averages and standard deviations of each test in both settings are shown in Appendix 0.B. Missing columns in the figures indicate the corresponding execution runs into error. A complete list of erroneous executions and explanations is listed in Appendix 0.A.
For the SPEC benchmark, on the x86 machine, S2malloc introduces 11.5% run-time overhead, smaller than two allocators – SlimGuard and DieHarder, and introduces the least memory overhead at 37.4%. On the AARCH machine, S2malloc introduces similar 15.5% run-time overhead, larger than the SlimGuard due to the fact that SlimGuard fails to run tests with frequent heap memory management operations. Running the PARSEC benchmark gives similar results, with smaller memory and run-time overheads.
We observe S2malloc and other memory allocators introduce larger overheads for tests with frequent heap memory management operations, for example, “ded” in PARSEC and “620” in SPEC. We investigate the costs of running malloc() and free() in the following section.
Run-time
Overhead
Memory
Overhead
S2malloc 189% 343%
SlimGuard 298% 250%
DieHarder 229% 92%
Guarder 56% 980%
Table 5: Normalized run-time and memory overheads of running mimalloc-benchmark
$\Delta$ Memory
Overhead
$\Delta$ Run-time
Overhead
0 Nearby -0.05% -0.42%
4 Nearby -0.13% +0.4%
4B Random N -12.35% -0.33%
12B Random N +68.23% +0.73%
50% Entropy +9.57% +0.49%
Table 6: Normalized memory and run-time Overhead changes compared with the default Settings.
6.2 Micro benchmarks
Figure 5: Execution time of glibc-simple
Figure 6: Run-time and memory overheads of running PARSEC with multi-threads
To further understand the overheads introduced by S2malloc, we investigate its performance using mimalloc-bench [9], composed of real-world and calibrated programs that allocate and free heap memory frequently. The results are shown in Table 6. Individual results are reported in Appendix 0.C. Running three tests with SlimGuard never returns (marked as ”-”) in Appendix 0.C. They are excluded from computing the SlimGuard average overheads.
All secure memory allocators incur larger overheads compared to running real-world (see § 6.3) or general-purpose benchmarks as mimalloc-bench tests operate the heap memory in a biased frequent way and some tests (e.g., ”leanN” generates the largest run-time overhead with S2malloc) counts the CPU ticks instead of seconds of finishing each call.
We take one test, glibc-simple from mimalloc-bench, from the glibc micro-benchmark suite [1] to further investigate the delays incurred in the two most common heap object management functions – malloc() and free(). The test times the execution of allocating and freeing a large number of blocks of a given size. We modify the benchmark and monitor the execution time of calling malloc() and free() separately. To investigate the time consumption of different sizes, we vary the block size S to be 16B, 128B, and 1KB, and change the number of allocated blocks N correspondingly, so that the total allocation size (N * S) is always 1000 MB. Results are presented in Figure 6, and is the average of 100 runs.
Generally, S2malloc takes more time to execute malloc() than all other compared memory allocators, and takes less time to execute free() than DieHarder but longer time than Guarder and SlimGuard. However, a significant overhead comes with our cryptographically secure canary implementation, which should be a standard adopted by all memory allocators. Although using hardware acceleration, the canary value is computed in each malloc() and free() calls, introducing a nonnegligible computation tax. After disabling this feature and using a fixed value as the canary, following the implementation of Guarder, although the execution time of both calls is still longer than Guarder and SlimGuard, it is comparable to others and the increased overhead is expected as S2malloc introduces extra security guarantees. For example, in 16B malloc call, S2malloc is 31% slower than SlimGuard and is 26% slower than SlimGuard in 16B free.
6.3 Performance on real world programs
To evaluate the performance of S2malloc in real-world environments, we run two servers: Nginx (1.18.0), and Lighttpd (1.4.71), and two databases: Redis (7.2.1), and SQLite (3.25.2) on the x86 machine. We use ApacheBench (ab) [37] 2.3 to test the throughput and delays using the Nginx default root page, of 613 bytes, as the requested page with 500 concurrent requests. On Redis, we use the same settings as its performance is measured in mimalloc-bench [9]. We use sqlite-bench [2] to measure the performance of SQLite. We report the results of performing random read and write operations in Table 7 and Table 8. We observe that applying S2malloc on these programs results in minimal throughput influence (even better throughput on Nginx and Redis). Running S2malloc delays the request response time on Nginx but not on Lighttpd. Applying all secure memory allocators increases memory consumption.
Nginx Lighttpd
Throughput Memory p50 Throughput Memory p50
S2malloc 9705.369 7867.2 56 11050.830 9425.6 44
Guarder 9496.210 13724 52 11146.580 11088.8 44
SlimGuard 6159.014 4935.2 81 11153.358 6428.0 44
DieHarder 8769.120 7396.8 57 11128.114 13626.4 44
Glibc 9564.754 3400.0 52 10742.708 5069.6 44
Table 7: Memory consumption (KB), throughput (requests/second), and delays (msec) for servers
Redis SQLite-Read SQLite-Write
Throughput Memory Throughput Memory Throughput Memory
S2malloc 218460.294 44688.8 771247.8791 17234.4 25146.73118 208592.0
Guarder 221245.016 46548.8 790263.9482 17451.2 25735.65366 193408.8
SlimGuard 219986.106 47397.6 766518.4731 14921.6 24456.21604 137659.2
DieHarder 221733.848 52419.2 781983.1092 19115.2 24613.93050 377160.8
Glibc 218155.764 50908.8 796812.7490 5914.4 25439.72566 132984.8
Table 8: Memory Consumption (KB) and throughput (requests/second) on databases
6.4 Performance with multi-threading
We run the memory allocators on the PARSEC benchmark with 4 and 16 threads separately using the x86 machine. We exclude the test “ray” from analysis as it cannot be executed with multiple threads and “vip“ as running it using Guarder with 16 threads causes a segmentation fault. Results are reported in Figure 6.
We observe that as the number of threads increases, S2malloc gradually introduces more run-time and memory overheads, as we use atomic instructions and maintain per-thread metadata. SlimGuard and DieHarder use single global metadata, and use lock to achieve multi-thread compatibility. While increasing the number of threads does not introduce extra memory overheads on the one hand, lock introduces more run-time overheads on the other hand. Guarder uses per-thread metadata but fails to use atomic instructions to update the metadata, causing racing conditions if multiple threads are handling adjacent blocks.
6.5 Influence with different parameters
In addition to the default settings, we also measure how different parameters, namely, nearby checking range, random offset entropy, and RIO entropy, influence the run-time and memory overheads. For the nearby checking range, we take 0 and 4 blocks as a comparison to the default setting: 2. For random allocation entropy, we take 4 bytes and 12 bytes as a comparison to the default 8 bytes. For random offset entropy, we reserve 50% of block size for RIO compared to the default 25%. Table 6 shows how different parameters influence the overheads.
We observe that changing the nearby checking range does not introduce observable differences for the memory overhead. The introduced delta is possibly due to server fluctuations. Using a larger nearby checking range introduces a larger run-time overhead, as S2malloc needs to compute and check more canary values. Using a larger random allocation entropy or RIO introduces both larger memory and run-time overheads.
7 Discussion
[…]
Limited UAF-read protection.Limited UAF-read protectionLimited UAF-read protection. S2malloc provides protection toward UAF-read based on the assumption that the attacker cannot distinguish the memory content stored in the victim data field from the content stored in other data fields. However, in the scenario where the victim object contains a field that the attacker could craft, e.g., a marker value like 0xdeadbeef, the victim field can be located by locating the crafted data field via UAF read(s).
This protection would be strong enough if the attack is complex enough that requires multiple successful UAF reads. However, it should be noted that as long as the dangling pointer can be used to access the memory, the attacker can read the data, and has the chance to identify the offset, which is the fundamental limitation of all statistically secure memory allocators. For a more comprehensive level of protection, other measures need to be taken. Complete mitigation of UAF-read attack is only achieved currently by either removing dangling pointers or pointer nullification.
[…]
Protecting kernel space memory.Protecting kernel space memoryProtecting kernel space memory. While S2malloc is a user-space memory allocator, its design is versatile enough to be adopted for protecting the kernel space. The Linux kernel operates with a limited number of heap objects, each field of which is equipped with finely calibrated permissions. Consequently, exploiting the kernel typically involves a series of intricate steps, aligns favorably with the design of S2malloc. Should the kernel adopts S2malloc into its structure, an attacker would be faced with the daunting task of achieving success in each individual attack phase, making a successful attack highly unlikely. We consider this as future work.
[…]
Increasing canary check frequency.Increasing canary check frequencyIncreasing canary check frequency. As the core of detecting invalid UAF writes, S2malloc checks the nearby canaries when a block is freed. While this design could virtually cover all memory slots for a program with frequent heap operations, UAF attack attempts cannot be detected if there is no malloc call that triggers S2malloc to check FBC.
While FBC checks could be integrated into other heap management functions call logics (e.g., free), the attacker can still possibly bypass the checking logic by never calling these functions. FBC checks need to be integrated into the regular program operations to fully address this issue, while imposing a significant performance overhead.
This problem, then, presents an opportunity for future research. A potential area of study could involve the strategic placement of canary checks at selected memory read locations. Additionally, an additional direction could be marking selected memory operations as sensitive that need to be integrated with FBC checks.
8 Conclusion
While statistically effective against all common heap exploitation techniques, state-of-the-art entropy-based heap allocators are not tailored to the active detection of unsuccessful exploitation attempts. As a result, in reality, to beat a randomization-based moving-target scheme, an attacker can simply launch the same attack, repeatedly, potentially with heap spraying, until success.
In this paper, we present S2malloc to fill the gap of exploitation attempt detection without compromising security and performance. In particular, we introduce three novel primitives to the design space of heap allocators: random in-block offset (RIO), free block canaries (FBC), and random bag layout (RBL). Combined with conventional BIBOP-style random allocation and heap canaries, S2malloc is able to maintain at least the same level of protection against other heap exploitations (e.g., overflows) and yet still achieves 69% and 96% protection rate in two attack scenarios, respectively, against UAF exploitation attempts targeting a 64 bytes object, while only incurs marginal performance overhead, making S2malloc practical to even production systems.
9 Acknowledgment
This work is funded in part by NSERC (RGPIN-2022-03325), the David R. Cheriton endowment, and gifts from Facebook, Amazon, Sui Foundation, and Intel Labs. We are grateful for the insightful feedback and guidance from our shepherd, Dr. R. Sekar which helped us to improve the paper significantly.
Appendix 0.A List of failed tests and corresponding exceptions
[…]
PARSEC x86.PARSEC x86PARSEC x86.
  • SlimGuard: ray(mmap error), flu(mmap error), fer(SIGSEGV), ded (SIGSEGV)
[…]
SPEC x86.SPEC x86SPEC x86.
  • SlimGuard: 600 (false positive double free), 602(Seg Fault), 623(Seg Fault), 631(Seg Fault), 644(Seg Fault), 657(Seg Fault)
  • DieHarder: 602(time out), 657(time out)
[…]
PARSEC AARCH.PARSEC AARCHPARSEC AARCH.
  • SlimGuard: ray(mmap error), flu(mmap error), fer(SIGSEGV), ded (SIGSEGV)
[…]
SPEC AARCH.SPEC AARCHSPEC AARCH.
  • SlimGuard: 600 (false positive double free), 602(Seg Fault), 623(Seg Fault), 631(Seg Fault), 644(Seg Fault), 657(Seg Fault)
  • DieHarder: All benchmarks (Too many open files error)
While running SPEC with DieHarder, test 602 runs unacceptably long, and possibly never returns. We killed the execution after running 1.5 hours, and we also mark it as invalid. As a comparison, running 602 on the AARCH machine with the default memory allocator only takes about 350 seconds.
NOTE: a PARSEC test is referenced by the first three letters of its name.
Appendix 0.B Run time and memory overheads of each PARSEC and SPEC test
Setup Run-time overhead Memory overhead
PARSEC
x86
SPEC
x86
PARSEC
AARCH
SPEC
AARCH
Table 9: Average and standard deviation of run time and memory overhead on PARSEC and SPEC benchmarks (x86 and AARCH).
Appendix 0.C Run time and memory overheads of each mimalloc-bench test
S2malloc SlimGuard DieHarder Guarder
cfrac 3.103 2.939 3.326 1.848
espresso 2.223 6.935 2.073 1.293
barnes 0.999 1.000 1.024 0.998
leanN 2.194 - 3.297 1.498
larsonN 9.313 23.250 7.033 1.748
mstressN 3.500 4.333 2.367 1.833
rptestN 5.383 18.131 3.243 1.212
gs 1.599 1.616 1.333 1.184
lua 1.579 2.364 1.337 1.200
alloc-test1 3.612 3.103 4.314 1.783
sh6benchN 5.392 - 40.675 2.228
sh8benchN 8.384 - 26.302 3.038
xmalloc-testN 2.640 4.052 5.136 1.519
cache-scratch1 1.000 1.000 1.007 1.000
glibc-simple 3.370 1244.895 2.801 1.824
glibc-thread 7.008 14.153 4.279 2.980
redis 1.006 1.004 1.008 0.999
average 2.891 3.981 3.298 1.561
Table 10: Normalized runtime overheads of mimalloc-bench
S2malloc SlimGuard DieHarder Guarder
cfrac 2.008 1.320 2.539 85.164
espresso 23.822 386.400 3.480 99.456
barnes 1.034 1.014 1.115 1.045
leanN 1.865 - 1.900 1.988
larsonN 11.189 5.667 0.986 17.966
mstressN 4.230 3.853 2.098 2.750
rptestN 12.056 5.119 3.224 5.661
gs 1.548 3.218 1.748 1.683
lua 1.760 1.353 1.333 1.494
alloc-test1 2.156 1.481 1.511 28.857
sh6benchN 1.103 - 1.697 3.109
sh8benchN 3.430 - 1.043 8.004
xmalloc-testN 580.270 8.197 3.032 1238.115
cache-scratch1 1.461 1.250 1.564 1.422
glibc-simple 3.592 239.045 3.778 260.030
glibc-thread 15.690 3.237 3.823 75.580
redis 1.231 1.099 1.249 1.203
average 4.340 4.502 1.922 10.800
Table 11: Normalized memory overheads of mimalloc-bench
Appendix 0.D Adapted code snippets to illustrate CVE-2015-6835 and its exploits
1HashTable *SESSION_VARS;
2
3PS_SERIALIZER_DECODE_FUNC(char* p, char *endptr) {
4  char *cursor = p;
5  hash_table_t *ht = INIT_HASHTABLE();
6  while (cursor < endptr) {
7    // for each item in the serialized stream
8    zval *name = PARSE_NEXT_ZVAL(&cursor);
9
10    zval *parsed = php_var_unserialize(&cursor, ht, endptr);
11    zval *stored = emalloc(sizeof(zval));
12    COPY_ZVAL(parsed, stored);
13    SESSION_VARS[name->value.str.val] = stored; // refcount = 1
14
15    zval_dtor(name);
16    zval_dtor(parsed);
17  }
18  free(ht);
19}
20zval* php_var_unserialize(char** pp, hash_table_t* ht, char* endptr) {
21  char *cursor = *pp;
22  zval *rval = emalloc(sizeof(zval));
23  INIT_ZVAL(rval); // refcount = 1
24
25  while (cursor < endptr) {
26    // for each element in the serialized item
27    zval *tmp;
28    if (*cursor == R) {
29        // reference to a previously parsed zval
30        cursor++;
31        tmp = PARSE_NEXT_ZVAL(&cursor);
32        tmp = ht[HASH_ZVAL(tmp)];
33    } else {
34        // parsing a new zval from the cursor
35        tmp = PARSE_NEXT_ZVAL(&cursor);
36    }
37    rval->value.ht[HASH_ZVAL(tmp)] = tmp;
38    zval_dtor(tmp);
39  }
40
41  *pp = cur;
42  ht[HASH_ZVAL(rval)] = rval;
43  return rval;
44}
45void zval_dtor(zval *p) {
46    p->refcount--;
47    if (p->refcount == 0) { free(p); }
48}
49
50/* ---------------------------------------------------------------- */
51
52int zend_echo_handler(char *name) {
53  zval *obj = SESSION_VARS[name];
54  obj->refcount++;
55  zend_print_variable(obj); // UAF read
56}
Figure 7: Adapted code snippets to illustrate CVE-2015-6835 and its exploits.
Appendix 0.E On The Formal Modeling of Probabilistic Use-After-Free Detection
To mathematically model how S2malloc provides defense against UAF, we make the following assumptions for the attacker and target program (which are consistent with our adversary model in § 3.1):
  • 1
    The goal of the attacker is to modify a sensitive field (e.g., a function pointer or an isadmin flag) in a specific type of object, a.k.a., a victim object, via memory writes over a dangling pointer (i.e., UAF-writes).
  • 2
    The attacker can obtain a dangling pointer through a bug in the program at any point of time during execution.
  • 3
    The program repetitively allocates and frees the type of objects targeted by the attacker (i.e., victim objects) during its execution. However, we do not assume that each victim object is freed before the next victim is allocated.
  • 4
    The attacker can either indirectly monitor or directly control the allocations of victim objects, i.e., the attacker knows when a victim object is allocated, but does not know the address of the allocation.
  • 5
    Any memory writes through the dangling pointer is conducted after the victim object is allocated.
  • 6
    If the intended sensitive field of a victim object is overridden, the attack succeeds; otherwise, the program continues to execute, allowing the attacker to repeat the exploitation effort unless detected by S2malloc (condition 7).
  • 7
    S2malloc checks FBCs on each heap allocation and detects the attack if any FBC is modified.
To simplify the illustration, we assume that the above execution logic is the only code logic that involves heap management. In real-world settings, attackers usually have an even lower success rate as memory slots can be allocated to other objects, which gives S2malloc more chances to check FBCs and detect UAF attempts.
[…]
Notation.NotationNotation. We denote the victim object size as $s$ which will be placed in a block of size $b$. Within a victim object, the sensitive data starts at the $s_{\_}1$ byte, and with length $l$. The RIO entropy is $e$ and obviously, $b\geq e+s$. The length of the FBC is $c$. The block-level entropy bit is $n$, i.e., each allocation of the victim object will fall in one of $r=2^{n}$ blocks.
0.E.1 Success rate of attack and defense per single attempt
In a block hosting a victim object, the first byte of the sensitive field is in the interval $[s_{\_}1,b-s+s_{\_}1)$. A reasonable attacker will always try to modify $l$ bytes of data starting at some byte within the interval. A smarter attacker will further leverage the knowledge that memory allocations are 16 bytes aligned (a convention from glibc). This implies that the RIO of a block is randomly chosen from one out of $1+\lfloor\dfrac{b-s}{16}\rfloor$ positions. Thus, if the attacker attempts to write $l$ bytes through the dangling pointer with a randomly guessed RIO value, the chance of success per trial is:
$A=\dfrac{1}{r}\vartriangleright\text{\small{the correct block}}\vartriangleleft\times\dfrac{1}{1+\lfloor\dfrac{b-s}{16}\rfloor}\vartriangleright\text{\small{the correct in-block RIO value}}\vartriangleleft=\dfrac{1}{r(1+\lfloor\dfrac{b-s}{16}\rfloor)}$ (1)
S2malloc puts an FBC randomly to any $c$ consecutive bytes in the block with the same probability. Hence, the probability that an FBC is modified by an $l$-byte write in the same block is reduced to the probability of selecting one $l$-byte chunk and one $c$-byte chunk randomly from a $b$-byte block and the two chunks overlaps by at least one byte.
$D$ $=\dfrac{2\sum_{\_}{i=0}^{c-2}(l+i)+\sum_{\_}{\_={c-1}}^{b-(l+c-1)}(l+c-1)\vartriangleright\text{\small{number of overlaps}}\vartriangleleft}{(b-l+1)(b-c+1)\vartriangleright\text{\small{number of ways to place $l$-byte and$c$-byte}}\vartriangleleft}$ (2)
$=\dfrac{b(l+c-1)-(l-1)^{2}-(c-1)^{2}-cl+1}{(b-l+1)(b-c+1)}$
The above equation holds when $b\geq l+2(c-1)$, which represents the most practical cases (i.e., lengths of both sensitive field and FBC are small) and favors the attacker. In fact, if either $l$ or $c$ is large enough relative to $b$, any $l$-byte write to the block will almost always corrupt the FBC and can be detected by S2malloc.
0.E.2 Strategy S1: repetitive UAF-writes to the same address
In this strategy, the attacker first obtains a dangling pointer ( 2) and holds the pointer for arbitrarily long. Every time the attacker notices a victim object is allocated ( 4), an $l$-byte UAF-write at the same offset through the same dangling pointer is conducted ( 5). This is essentially repetitive UAF-writes to the same address.
As S2malloc only detects UAF attempts when a victim object is allocated, we use round $i$ to represent the $i$-th allocation of a victim object after the attacker obtains the dangling pointer and conducts the UAF-write. Effectively, after round $i$, S2malloc should have checked FBCs $i$ times to catch the UAF attempt.
We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches round $i$. By definition, _$\mathbb{P}_{\_}{e}^{1}=1-A$, i.e., when the attacker’s first UAF-write is not successful in achieving the goal ( 1). Suppose the execution has reached round $i$, the probability that the repetitive UAF-writes is detected at this particular round is
[…]
Based on this, we can derive the inductive definition for _$\mathbb{P}_{\_}{e}^{i}$:
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ reaches round $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\dfrac{2d+1}{r}\times D)\vartriangleright\text{\small{undetected}}\vartriangleleft$ (4)
[…]
Limiting program execution to an upper bound of $K$ rounds, the chance of attacker and S2malloc wins, respectively, is:
_$\mathbb{P}_{\_}{attack}^{K}=A+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times A)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\dfrac{2d+1}{r}\times D)$ (5)
0.E.3 Strategy S2: UAF-writes through fresh dangling pointers
Unlike § 0.E.2, the attacker does not hold a dangling pointer indefinitely, instead, the attacker obtains a fresh dangling pointer ( 2) if a prior UAF-write attempt is not successful. After obtaining a dangling pointer, if the attacker notices a victim object is allocated ( 4), an $l$-byte UAF-write through the fresh dangling pointer is conducted ( 5). This essentially means that every UAF-write is likely on a different address.
More importantly, as S2malloc only detects UAF attempts when a victim object is allocated, this strategy effectively creates a turn-based game between the attacker and S2malloc where in each round, the attacker makes the move of obtaining a dangling pointer and conducting a UAF-write while S2malloc makes the move of checking FBCs and allocating a new victim object (if FBCs checked are intact). The game ends when either the attacker or S2malloc wins.
We use round $i$ to represent the $i$-th round of the game. In each round, the attacker makes the first move and S2malloc follows. We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches S2malloc’s turn in round $i$ (to be consistent with the notation in § 0.E.2). By definition, _$\mathbb{P}_{\_}{e}^{1}=1-A$, i.e., the attacker’s first UAF-write is not successful in achieving the goal ( 1).
To calculate the detection rate by S2malloc, we rephrase the question to a classical combinatorics question: there exists $r$ balls in a box where each time the attacker picks one ball randomly (i.e., the block referred by the dangling pointer), colors it with probability $D$ (i.e., corrupts the FBC in the block), and puts the ball back to the box. A ball cannot be uncolored once it is colored (because the attacker does not undo a UAF-write). We use _$\mathbb{Q}_{\_}i$ to denote the probability that an arbitrary ball in the box is not colored (i.e., a block with its FBCs integral) after $i$ rounds.
_$\mathbb{Q}_{\_}i=(\dfrac{r-1}{r}\vartriangleright\text{\small{ball not selected}}\vartriangleleft+\dfrac{1}{r}\times(1-D)\vartriangleright\text{\small{ball selected but not colored}}\vartriangleleft)^{i}=(\dfrac{r-D}{r})^{i}$ (6)
Therefore, at round $i$, there will be, by expected value, $r\cdot Q_{\_}i$ balls remain uncolored in the box. The detection rate of S2malloc at round $i$ will be the same as the probability of selecting $12d+$ 1 consecutive balls from a string of $r$ balls where at least one of the selected balls is colored. The detection rate is denoted as _$\mathbb{P}_{\_}{d}^{i}$ and calculated as:
_$\mathbb{P}_{\_}{d}^{i}=1-\dfrac{Q_{\_}ir-2d}{r-2d}\cdot\prod_{\_}{i=0}^{2d}(\dfrac{Q_{\_}ir-i}{r-i})$ (7)
Based on this, we can derive the inductive definition for _$\mathbb{P}_{\_}{e}^{i}$:
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\mathbb{P}_{\_}{d}^{i})\vartriangleright\text{\small{undetected}}\vartriangleleft\times$ (8)
[…]
Limiting program execution to an upper bound of $K$ rounds, the chance of attacker and S2malloc wins, respectively, is:
_$\mathbb{P}_{\_}{attack}^{K}=A+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times A)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\mathbb{P}_{\_}{d}^{i})$ (9)
0.E.4 Strategy S1-spray: repetitive UAF-writes to the same address with spraying
This strategy operates similarly to the strategy in § 0.E.2: the attacker first obtains a dangling pointer ( 2) and holds the pointer for arbitrarily long. However, generalized from § 0.E.2 in which the attacker conducts a UAF-write after one allocation of a victim object ( 4), the attacker waits until there are $m$ victim objects newly allocated and alive (i.e., not freed yet) and then issues the UAF-write. Effectively, the attacker is trying to diligently spray the heap by victim objects to increase its chance of success. The UAF-write is still an $l$-byte memory write at the same offset through the same dangling pointe ( 5). This is essentially repetitive UAF-writes to the same address, similar to § 0.E.2.
Consistent with the analysis in § 0.E.2, we still use round $i$ to represent the $i$-th allocation of a victim object after the attacker obtains the dangling pointer and conducts the UAF-write. Effectively, after round $i$, S2malloc should have checked FBCs $i$ times to catch the UAF attempt.
We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches round $i$. By definition, _$\mathbb{P}_{\_}{e}^{1}=1-mA$, i.e., when the attacker’s first UAF-write is not successful in achieving the goal ( 1). Note that the attacker success rate increases as there are $m$ victim objects alive and the attack succeeds as long as the sensitive field in any one of them is overridden by the UAF-write—this is essentially the advantage of heap spraying.
And yet, consistent with § 0.E.2, this UAF-write can corrupt one FBC at most. Hence, suppose the execution has reached round $i$, the probability that the repetitive UAF-writes is detected at this particular round is still Equation 3. With attacker’s success rate improved, the inductive definition for _$\mathbb{P}_{\_}{e}^{i}$ in this strategy will be:
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\dfrac{2d+1}{r}\times D)\vartriangleright\text{\small{undetected}}\vartriangleleft\times$ (10)
[…]
Limiting program execution to an upper bound of $K$ rounds, the chance of attacker and S2malloc wins, respectively, is:
_$\mathbb{P}_{\_}{attack}^{K}=mA+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times mA)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\dfrac{2d+1}{r}\times D)$ (11)
0.E.5 Strategy S2-spray: UAF-writes through fresh dangling pointers with spraying
In this strategy, the attacker still sprays the heap such that overriding the sensitive field in any the $m$ victim objects achieves the goal (like the strategy in § 0.E.4). Similar to § 0.E.3, the attacker does not hold a dangling pointer indefinitely, instead, the attacker obtains a fresh dangling pointer ( 2) if a prior UAF-write attempt is not successful and use it to launch a UAF-write attack in the current round. In each round, the attacker makes the first move and S2malloc follows. We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches S2malloc’s turn in round $i$ (to be consistent with the notation in § 0.E.2§ 0.E.3, and § 0.E.4). By definition, _$\mathbb{P}_{\_}{e}^{1}=1-mA$, i.e., the attacker’s first UAF-write is not successful in achieving the goal ( 1) even after spraying $m$ victim objects.
Similar to § 0.E.3, we use _$\mathbb{Q}_{\_}i$ to denote the probability that an arbitrary ball in the box is not colored (i.e., a block with its FBCs integral) after $i$ rounds. _$\mathbb{Q}_{\_}i$ and _$\mathbb{P}_{\_}{d}^{i}$ can be computed using the same formula as in § 0.E.3
Based on this, we can derive the inductive definition for _$\mathbb{P}_{\_}{e}^{i}$:
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\mathbb{P}_{\_}{d}^{i})\vartriangleright\text{\small{undetected}}\vartriangleleft\times$ (12)
[…]
Limiting program execution to an upper bound of $K$ rounds, the chance of attacker and S2malloc wins, respectively, is:
_$\mathbb{P}_{\_}{attack}^{K}=mA+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times mA)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\mathbb{P}_{\_}{d}^{i})$ (13)